OpenStudy (anonymous):

\[Evaluate ==> \sin ^{2}xcosxdx\]

OpenStudy (lgbasallote):

do you mean integral?

OpenStudy (anonymous):

You can write: \[\sin^2xcosx = (1-\cos^2x).(cosx) = cosx - \cos^3x\] So. \[\int\limits cosx.dx - \int\limits \cos^2x dx\]

OpenStudy (anonymous):

Try to solve this..

OpenStudy (lgbasallote):

careful

OpenStudy (anonymous):

You are scaring me..

OpenStudy (lgbasallote):

you can just simply do \[\int \sin^2 x \cos x dx\] let u = \(\sin x)\) du = \(\cos x\) \[\int \sin^2 x \cos x dx \implies \int u^2 du\] now it's just power rule

OpenStudy (lgbasallote):

no need to do the complex integral such as cos^2 x dx

OpenStudy (anonymous):

oh yeah..

OpenStudy (lgbasallote):

wait...how do you integrate cos^3 dx anyway o.O

OpenStudy (lgbasallote):

you'll have to go back to sin^2 x cos x lol

OpenStudy (anonymous):

haha. :D

OpenStudy (anonymous):

NO..

OpenStudy (anonymous):

\[\cos^3x = \frac{\cos(3x) + 3\cos(x)}{4}\]